OpenStudy (anonymous):

find the multiplicative inverse of 3/2? explain.

OpenStudy (anonymous):

is it -3/2?

TheSmartOne (thesmartone):

The multiplicative inverse of a fraction a/b is b/a.

TheSmartOne (thesmartone):

So what do you think is the inverse now?

TheSmartOne (thesmartone):

@mcnabkt

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

im still confused

TheSmartOne (thesmartone):

\[\frac{ a }{ b }=\frac{ 3 }{ 2 }\] \[\frac{ b }{ a }=\frac{ ? }{ ? }\]

TheSmartOne (thesmartone):

Just switch the numbers

OpenStudy (anonymous):

ohh 2/3

TheSmartOne (thesmartone):

\[\frac{ a }{ b } \times \frac{ b }{ a }=1\] and \[\frac{ 3 }{ 2 }\times \frac{ ? }{ ? }=1\]

OpenStudy (anonymous):

1/2?

TheSmartOne (thesmartone):

The number and its multiplicative inverse always multiply together to equal 1

TheSmartOne (thesmartone):

\[\frac{ 3 }{ 2 }\times \frac{ 2 }{ 3 }=1\]

OpenStudy (anonymous):

okay well what about the multiplicative inverse for 1/13? would it be 13/1?

TheSmartOne (thesmartone):

Exactly.. Also remember any number is the same thing as over one

TheSmartOne (thesmartone):

For example\[5=\frac{ 5 }{ 1 }\]

OpenStudy (anonymous):

so i put 13/1 for the answer?

TheSmartOne (thesmartone):

Or you can just put 13

OpenStudy (anonymous):

oh i can lol

OpenStudy (anonymous):

uhmm 4/17 would be.... 2?

TheSmartOne (thesmartone):

17/4 not 2

TheSmartOne (thesmartone):

remember the reciprocal of the number is its multiplicative inverse

OpenStudy (anonymous):

ughh well you just said or you can.. nvm different problem LOL so i just switch the numerator and denominator?

TheSmartOne (thesmartone):

yes

TheSmartOne (thesmartone):

that is what a reciprocal is

OpenStudy (anonymous):

ohh okay well i need help with a couple more

OpenStudy (anonymous):

what about multiplicative inverse of a fraction? wait were we just going over this topic?

TheSmartOne (thesmartone):

yes

OpenStudy (anonymous):

oh okay see i told you its very hard of me understanding lol

TheSmartOne (thesmartone):

For the multiplicative inverse of a number such as 5 imagine it as 5/1 and then its multiplicative inverse is 1/5

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

hold on i got a few more

OpenStudy (anonymous):

i got 10 out of 10

TheSmartOne (thesmartone):

Hooray.

OpenStudy (anonymous):

lol i was surprised

OpenStudy (anonymous):

do you understand product of fractions?

TheSmartOne (thesmartone):

yes

OpenStudy (anonymous):

okay good cause thats what im doing now

OpenStudy (anonymous):

how are you so smart? and what GPA do you have lol

OpenStudy (anonymous):

okay so 2/7 x 1/3 would be 2/21?

TheSmartOne (thesmartone):

Yup..

TheSmartOne (thesmartone):

multiplication in fractions\[\frac{ a }{ b }\times \frac{ x }{ y }=\frac{ ax }{ by }\]

OpenStudy (anonymous):

what about -1/4x (-3/7)? would it be -3/28?

TheSmartOne (thesmartone):

\[\frac{ -1 }{ 4 }\times \frac{ -3 }{7 }=\frac{ -1\times-3 }{ 4\times7 }=\frac{ 3 }{ 28 }\]

OpenStudy (anonymous):

yay i was right

TheSmartOne (thesmartone):

No you put -3/28 not 3/28

OpenStudy (anonymous):

what about 2 1/4 x 1 3/8? i got 2 3/32 and it was wrong and i put 3/28 and got it right lol idk my class thing is a little weird it says its wrong when its right and right when its wrong

TheSmartOne (thesmartone):

\[2 \frac{ 1 }{ 4}=\frac{ 9 }{ 4 }\] \[1 \frac{ 3 }{ 8}=\frac{ 11 }{ 8 }\] \[\frac{ 9 }{ 4 }\times \frac{ 11 }{ 8 }=\frac{ 9\times11 }{ 4 \times8 }=\frac{ 99 }{ 32 }\]

OpenStudy (anonymous):

ohh i guess i forgot to work it out

OpenStudy (anonymous):

i dont usually work my problems out like that but im supposed to

OpenStudy (anonymous):

- 1/6 x (- 3/4) how is -3/24 wrong?

TheSmartOne (thesmartone):

\[\frac{ -1 }{ 6 }\times \frac{ -3 }{ 4 }=\frac{ -1\times-3 }{ 6\times4 }=\frac{ 3 }{ 24 }=\frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

oh i forgot to simplify i guess

TheSmartOne (thesmartone):

Also the - sign

TheSmartOne (thesmartone):

\[\frac{ 3 }{ 24}\neq \frac{ -3 }{ 24 }\]

OpenStudy (anonymous):

-2 2/3 x 4 2/5 would be... -8 4/15? then i simplify it and i get